If 5kg of body falls from the 30 m of height and if all its mechanical energy is converted into heat produced will be?

Grade:12th pass

1 Answers

Shùbháñk Dîxít
7 years ago
m=5 kg ,h=30 since ball is dropped therefore u=0 now v^2-u^2=2*g*hv^2-0^2=2*9.8*30 v^2=58.8 m/s since mechanical energy is converted into heat energy therefore mechanical energy = kinetic energy = heat energy heat energy = (1/2)*m*(v^2) =1/2*5*58.8= 147 joule